Given function
f(x) = 4 x^2 - 5 x^1 + 1 ,
find the critical numbers of the function.
Enter the largest non-zero one. (Keep 2 decimal places)

Answer by richard1234(7193)

f'(x) = 8x - 5

Critical point when 8x - 5 = 0, or x = 5/8 = 0.625.